### v1

@@ 9 Experiments
In the special case where we know that the images considered are captured using a set of cameras that have a similar lens, i.e., undistortion parameters lie in a small interval, solvers that sample undistortion parameters to estimate the pose can get a noticeable performance boost. We conducted two synthetic experiments on images from 12 scenes from the ETH3D dataset [41], in which we applied uniformly random distortion from the ranges (1) $[-1.7,-1.3]$, and (2) $[-0.9,-0.5]$. We refer to the former as Wide-angle lens, and to the latter as Medium-angle lens. The tested sampling solvers consider only one undistortion parameter, in particular the mean value of the undistortion interval in the corresponding setup. The results are reported in Tab. 9 and 10, for the Wide-angle lens, and Medium-angle lens setups, respectively. In both scenarios, we consider the case of equal unknown radial distortion (i.e., $\lambda_{1}=\lambda_{2}$). As can be seen, combining the 7pt $\mathbf{F}$ solver with a sampled radial undistortion parameter leads to a comparable accuracy as employing the minimal 8pt $\mathbf{F}\lambda$ solver at much faster run-times. At the same time, the combination of the sample and the 7pt $\mathbf{F}$ solver is more accurate and faster compared to using the non-minimal 9pt $\mathbf{F}\lambda$ solver.
Figure 9: Robustness to varying levels of equal distortion: We show the estimated radial undistortion parameter, the pose error, and the rotation and translation errors for the ETH3D dataset. An x-axis value of $-0.3$ means that we apply a distortion that can be undistorted with the undistortion parameter $-0.3$. The cyan line indicates the ground truth $\lambda$ value on the y-axis.
Figure 10: Robustness to varying levels of equal distortion: We show the estimated radial undistortion parameter, the pose error, and the rotation and translation errors for the Sacre Coeur scene from the Phototourism dataset. An x-axis value of $-0.3$ means that we apply a distortion that can be undistorted with the undistortion parameter $-0.3$. The cyan line indicates the ground truth $\lambda$ value on the y-axis.
@@
Fig. 9 extends the results shown in Fig. 1 of the main paper for 12 scenes from ETH3D dataset [41]. Fig. 1 in the main paper showed the estimated undistortion parameter and the pose error. In addition to these two error measures, Fig. 9 also shows the rotation and translation errors.
@@
Our explanation is that the "undistorted" images provided by the dataset are not fully undistorted, i.e., COLMAP [39] (which was used to estimate ground truth poses and intrinsic parameters) did not fully predict the distortion of each image. Rather, a small distortion remains. COLMAP uses Approach (1) discussed in the introduction of the main paper, i.e., COLMAP handles radial distortion only during the bundle adjustment stage of the reconstruction process. Our results seem to indicate that modeling distortions in earlier stages could be beneficial. Due to their simplicity, efficiency, and effectiveness, the sampling-based approaches proposed in our work are promising candidates for this task.
[…]
@@
Minimal Refinement Sample AVG $(^{\circ})$downarrow MED $(^{\circ})$downarrow AUC@5 $\uparrow$ @10 @20 AVG $\epsilon(\lambda)$downarrow MED $\epsilon(\lambda)$downarrow Time (ms) $\downarrow$
$\lambda_{1}=\lambda_{2}$ 7pt $\mathbf{F}$ $\mathbf{F}$ 0 37.97 15.97 0.15 0.24 0.36 0.87 0.87 1132.69
7pt $\mathbf{F}$ $\mathbf{F}\lambda$ 00 23.30 11.38 0.56 0.61 0.67 0.26 0.03 1125.28
7pt $\mathbf{F}$ $\mathbf{F}\lambda_{1}\lambda_{2}$ 00 25.40 11.79 0.54 0.59 0.65 0.35 0.07 1126.50
@@
$\lambda_{1}\neq\lambda_{2}$ 7pt $\mathbf{F}$ $\mathbf{F}$ 0 45.01 23.81 0.09 0.14 0.25 0.87 0.87 1147.94
7pt $\mathbf{F}$ $\mathbf{F}\lambda_{1}\lambda_{2}$ 00 29.03 13.43 0.47 0.53 0.59 0.38 0.12 1160.31
@@
Minimal Refinement Sample AVG $(^{\circ})$downarrow MED $(^{\circ})$downarrow AUC@5 $\uparrow$ @10 @20 AVG $\epsilon(\lambda)$downarrow MED $\epsilon(\lambda)$downarrow Time (ms) $\downarrow$
$\lambda_{1}=\lambda_{2}$ 7pt $\mathbf{F}$ $\mathbf{F}$ 0 20.22 3.20 0.45 0.58 0.68 0.15 0.15 1121.63
7pt $\mathbf{F}$ $\mathbf{F}\lambda$ 00 17.34 0.70 0.66 0.71 0.76 0.07 0.01 $1$.57
7pt $\mathbf{F}$ $\mathbf{F}\lambda_{1}\lambda_{2}$ 00 19.20 0.87 0.64 0.70 0.74 0.17 0.03 1112.44
@@
$\lambda_{1}\neq\lambda_{2}$ 7pt $\mathbf{F}$ $\mathbf{F}$ 0 23.31 4.66 0.34 0.48 0.60 0.15 0.15 1125.48
7pt $\mathbf{F}$ $\mathbf{F}\lambda_{1}\lambda_{2}$ 00 19.86 1.00 0.62 0.68 0.73 0.18 0.03 1117.94
@@
Minimal Refinement Sample AVG $(^{\circ})$downarrow MED $(^{\circ})$downarrow AUC@5 $\uparrow$ @10 @20 AVG $\epsilon(\lambda)$downarrow MED $\epsilon(\lambda)$downarrow Time (ms) $\downarrow$
$\lambda_{1}=\lambda_{2}$ 7pt $\mathbf{F}$ $\mathbf{F}$ 0 44.12 20.78 0.07 0.13 0.26 1.10 1.09 1136.27
7pt $\mathbf{F}$ $\mathbf{F}\lambda$ $-0.9$ 18.63 10.84 0.63 0.69 0.73 0.12 0.02 1102.13
7pt $\mathbf{F}$ $\mathbf{F}\lambda_{1}\lambda_{2}$ $-0.9$ 20.80 11.17 0.60 0.66 0.71 0.24 0.05 1107.32
@@
$\lambda_{1}\neq\lambda_{2}$ 7pt $\mathbf{F}$ $\mathbf{F}$ 0 48.82 26.36 0.05 0.10 0.21 1.10 1.10 1144.53
7pt $\mathbf{F}$ $\mathbf{F}\lambda_{1}\lambda_{2}$ $-0.9$ 23.66 11.66 0.55 0.61 0.67 0.26 0.06 1124.89
@@
Minimal Non-Minimal Sample AVG $(^{\circ})$downarrow MED $(^{\circ})$downarrow AUC@5 $\uparrow$ @10 @20 AVG $\epsilon(\lambda)$downarrow MED $\epsilon(\lambda)$downarrow Time (ms) $\downarrow$
$\lambda_{1}=\lambda_{2}$ 7pt $\mathbf{F}$ 7pt $\mathbf{F}$ 0 13.11 5.72 0.21 0.38 0.55 0.86 0.85 190.04
@@
$\lambda_{1}\neq\lambda_{2}$ 7pt $\mathbf{F}$ 7pt $\mathbf{F}$ 0 15.14 7.04 0.15 0.32 0.50 0.86 0.86 140.78
@@
Minimal Non-Minimal Sample AVG $(^{\circ})$downarrow MED $(^{\circ})$downarrow AUC@5 $\uparrow$ @10 @20 AVG $\epsilon(\lambda)$downarrow MED $\epsilon(\lambda)$downarrow Time (ms) $\downarrow$
$\lambda_{1}=\lambda_{2}$ 7pt $\mathbf{F}$ 7pt $\mathbf{F}$ 0 30.22 22.80 0.05 0.11 0.23 0.85 0.84 170.37
@@
$\lambda_{1}\neq\lambda_{2}$ 7pt $\mathbf{F}$ 7pt $\mathbf{F}$ 0 31.36 25.39 0.03 0.09 0.20 0.86 0.86 141.03
@@
Minimal Non-Minimal Sample AVG $(^{\circ})$downarrow MED $(^{\circ})$downarrow AUC@5 $\uparrow$ @10 @20 AVG $\epsilon(\lambda)$downarrow MED $\epsilon(\lambda)$downarrow Time (ms) $\downarrow$
$\lambda_{1}=\lambda_{2}$ 7pt $\mathbf{F}$ 7pt $\mathbf{F}$ 0 30.74 22.21 0.06 0.15 0.26 0.87 0.87 180.49
@@
$\lambda_{1}\neq\lambda_{2}$ 7pt $\mathbf{F}$ 7pt $\mathbf{F}$ 0 31.60 23.63 0.04 0.11 0.23 0.86 0.87 160.39
@@
Minimal Non-Minimal Sample AVG $(^{\circ})$downarrow MED $(^{\circ})$downarrow AUC@5 $\uparrow$ @10 @20 AVG $\epsilon(\lambda)$downarrow MED $\epsilon(\lambda)$downarrow Time (ms) $\downarrow$
$\lambda_{1}=\lambda_{2}$ 7pt $\mathbf{F}$ 7pt $\mathbf{F}$ 0 17.24 6.45 0.21 0.36 0.51 0.86 0.85 160.49
@@
$\lambda_{1}\neq\lambda_{2}$ 7pt $\mathbf{F}$ 7pt $\mathbf{F}$ 0 18.78 8.04 0.16 0.31 0.47 0.86 0.86 161.19
@@
Minimal Non-Minimal Sample AVG $(^{\circ})$downarrow MED $(^{\circ})$downarrow AUC@5 $\uparrow$ @10 @20 AVG $\epsilon(\lambda)$downarrow MED $\epsilon(\lambda)$downarrow Time (ms) $\downarrow$
$\lambda_{1}=\lambda_{2}$ 7pt $\mathbf{F}$ 7pt $\mathbf{F}$ 0 29.09 21.46 0.05 0.12 0.24 0.86 0.84 150.03
@@
$\lambda_{1}\neq\lambda_{2}$ 7pt $\mathbf{F}$ 7pt $\mathbf{F}$ 0 31.74 24.68 0.03 0.09 0.20 0.86 0.85 150.01
@@
Minimal Non-Minimal Sample AVG $(^{\circ})$downarrow MED $(^{\circ})$downarrow AUC@5 $\uparrow$ @10 @20 AVG $\epsilon(\lambda)$downarrow MED $\epsilon(\lambda)$downarrow Time (ms) $\downarrow$
$\lambda_{1}=\lambda_{2}$ 7pt $\mathbf{F}$ 7pt $\mathbf{F}$ 0 36.74 31.81 0.05 0.10 0.19 0.86 0.85 140.01
@@
$\lambda_{1}\neq\lambda_{2}$ 7pt $\mathbf{F}$ 7pt $\mathbf{F}$ 0 39.80 35.68 0.02 0.06 0.13 0.86 0.86 140.71
@@
Minimal Non-Minimal Sample AVG $(^{\circ})$downarrow MED $(^{\circ})$downarrow AUC@5 $\uparrow$ @10 @20 AVG $\epsilon(\lambda)$downarrow MED $\epsilon(\lambda)$downarrow Time (ms) $\downarrow$
$\lambda_{1}=\lambda_{2}$ 7pt $\mathbf{F}$ 7pt $\mathbf{F}$ 0 25.16 16.80 0.08 0.17 0.31 0.88 0.88 1120.22
@@
$\lambda_{1}\neq\lambda_{2}$ 7pt $\mathbf{F}$ 7pt $\mathbf{F}$ 0 32.47 24.35 0.04 0.10 0.21 0.86 0.84 1203.38
@@
Minimal Non-Minimal Sample AVG $(^{\circ})$downarrow MED $(^{\circ})$downarrow AUC@5 $\uparrow$ @10 @20 AVG $\epsilon(\lambda)$downarrow MED $\epsilon(\lambda)$downarrow Time (ms) $\downarrow$
$\lambda_{1}=\lambda_{2}$ 7pt $\mathbf{F}$ 9pt $\mathbf{F}\lambda$ $-1.5$ 15.85 2.43 0.38 0.52 0.64 0.13 0.08 1170.05
7pt $\mathbf{F}$ 12pt $\mathbf{F}\lambda_{1}\lambda_{2}$ $-1.5$ 16.06 2.80 0.34 0.51 0.63 0.09 0.09 1180.12
@@
Minimal Non-Minimal Sample AVG $(^{\circ})$downarrow MED $(^{\circ})$downarrow AUC@5 $\uparrow$ @10 @20 AVG $\epsilon(\lambda)$downarrow MED $\epsilon(\lambda)$downarrow Time (ms) $\downarrow$
$\lambda_{1}=\lambda_{2}$ 7pt $\mathbf{F}$ 9pt $\mathbf{F}\lambda$ $-0.7$ 16.64 2.58 0.37 0.51 0.62 0.17 0.08 1205.99
7pt $\mathbf{F}$ 12pt $\mathbf{F}\lambda_{1}\lambda_{2}$ $-0.7$ 15.98 3.11 0.32 0.49 0.62 0.10 0.09 1220.47

### v2

@@ 9 Experiments
In the special case where we know that the images considered are captured using a set of cameras that have a similar lens, i.e., undistortion parameters lie in a small interval, solvers that sample undistortion parameters to estimate the pose can get a noticeable performance boost. We conducted two synthetic experiments on images from 12 scenes from the ETH3D dataset [Schops_2017_CVPR], in which we applied uniformly random distortion from the ranges (1) [1.7,1.3][-1.7,-1.3], and (2) [0.9,0.5][-0.9,-0.5]. We refer to the former as Wide-angle lens, and to the latter as Medium-angle lens. The tested sampling solvers consider only one undistortion parameter, in particular the mean value of the undistortion interval in the corresponding setup. The results are reported in Tab. 9 and 10, for the Wide-angle lens, and Medium-angle lens setups, respectively. In both scenarios, we consider the case of equal unknown radial distortion (i.e., $\lambda_{1}=\lambda_{2}$). As can be seen, combining the 7pt $\mathbf{F}$ solver with a sampled radial undistortion parameter leads to a comparable accuracy as employing the minimal 8pt $\mathbf{F}\lambda$ solver at much faster run-times. At the same time, the combination of the sample and the 7pt $\mathbf{F}$ solver is more accurate and faster compared to using the non-minimal 9pt $\mathbf{F}\lambda$ solver.
Figure 9: Robustness to varying levels of equal distortion: We show the estimated radial undistortion parameter, the pose error, and the rotation and translation errors for the ETH3D dataset. An x-axis value of 0.3-0.3 means that we apply a distortion that can be undistorted with the undistortion parameter 0.3-0.3. The cyan line indicates the ground truth $\lambda$ value on the y-axis.
Figure 10: Robustness to varying levels of equal distortion: We show the estimated radial undistortion parameter, the pose error, and the rotation and translation errors for the Sacre Coeur scene from the Phototourism dataset. An x-axis value of 0.3-0.3 means that we apply a distortion that can be undistorted with the undistortion parameter 0.3-0.3. The cyan line indicates the ground truth $\lambda$ value on the y-axis.
@@
Fig. 9 extends the results shown in Fig. 1 of the main paper for 12 scenes from ETH3D dataset [Schops_2017_CVPR]. Fig. 1 in the main paper showed the estimated undistortion parameter and the pose error. In addition to these two error measures, Fig. 9 also shows the rotation and translation errors.
@@
Our explanation is that the "undistorted" images provided by the dataset are not fully undistorted, i.e., COLMAP [Schoenberger2016CVPR] (which was used to estimate ground truth poses and intrinsic parameters) did not fully predict the distortion of each image. Rather, a small distortion remains. COLMAP uses Approach (1) discussed in the introduction of the main paper, i.e., COLMAP handles radial distortion only during the bundle adjustment stage of the reconstruction process. Our results seem to indicate that modeling distortions in earlier stages could be beneficial. Due to their simplicity, efficiency, and effectiveness, the sampling-based approaches proposed in our work are promising candidates for this task.
Table 5: Results for {($\{$(Scenario A) Wild, (Scenario B) Small distortion, (Scenario C) Visible distortion}\} on 12 scenes from ETH3D dataset, using the Poselib framework and SP+LG matches.
@@
Minimal Refinement Sample AVG $(^{\circ})$ $\downarrow$ MED $(^{\circ})$ $\downarrow$ AUC@5 $\uparrow$ @10 @20 AVG $\epsilon(\lambda)$ $\downarrow$ MED $\epsilon(\lambda)$ $\downarrow$ Time (ms) $\downarrow$
$\lambda_{1}=\lambda_{2}$
7pt $\mathbf{F}$ $\mathbf{F}$ 0 37.97 15.97 0.15 0.24 0.36 0.87 0.87 1132.69
7pt $\mathbf{F}$ $\mathbf{F}\lambda$ 0 23.30 11.38 0.56 0.61 0.67 0.26 0.03 1125.28
7pt $\mathbf{F}$ $\mathbf{F}\lambda_{1}\lambda_{2}$ 0 25.40 11.79 0.54 0.59 0.65 0.35 0.07 1126.50
@@
$\lambda_{1}\neq\lambda_{2}$
7pt $\mathbf{F}$ $\mathbf{F}$ 0 45.01 23.81 0.09 0.14 0.25 0.87 0.87 1147.94
7pt $\mathbf{F}$ $\mathbf{F}\lambda_{1}\lambda_{2}$ 0 29.03 13.43 0.47 0.53 0.59 0.38 0.12 1160.31
@@
Minimal Refinement Sample AVG $(^{\circ})$ $\downarrow$ MED $(^{\circ})$ $\downarrow$ AUC@5 $\uparrow$ @10 @20 AVG $\epsilon(\lambda)$ $\downarrow$ MED $\epsilon(\lambda)$ $\downarrow$ Time (ms) $\downarrow$
$\lambda_{1}=\lambda_{2}$
7pt $\mathbf{F}$ $\mathbf{F}$ 0 20.22 3.20 0.45 0.58 0.68 0.15 0.15 1121.63
7pt $\mathbf{F}$ $\mathbf{F}\lambda$ 0 17.34 0.70 0.66 0.71 0.76 0.07 0.01 1111.57
7pt $\mathbf{F}$ $\mathbf{F}\lambda_{1}\lambda_{2}$ 0 19.20 0.87 0.64 0.70 0.74 0.17 0.03 1112.44
@@
$\lambda_{1}\neq\lambda_{2}$
7pt $\mathbf{F}$ $\mathbf{F}$ 0 23.31 4.66 0.34 0.48 0.60 0.15 0.15 1125.48
7pt $\mathbf{F}$ $\mathbf{F}\lambda_{1}\lambda_{2}$ 0 19.86 1.00 0.62 0.68 0.73 0.18 0.03 1117.94
@@
Minimal Refinement Sample AVG $(^{\circ})$ $\downarrow$ MED $(^{\circ})$ $\downarrow$ AUC@5 $\uparrow$ @10 @20 AVG $\epsilon(\lambda)$ $\downarrow$ MED $\epsilon(\lambda)$ $\downarrow$ Time (ms) $\downarrow$
$\lambda_{1}=\lambda_{2}$
7pt $\mathbf{F}$ $\mathbf{F}$ 0 44.12 20.78 0.07 0.13 0.26 1.10 1.09 1136.27
7pt $\mathbf{F}$ $\mathbf{F}\lambda$ 0.9-0.9 18.63 10.84 0.63 0.69 0.73 0.12 0.02 1102.13
7pt $\mathbf{F}$ $\mathbf{F}\lambda_{1}\lambda_{2}$ 0.9-0.9 20.80 11.17 0.60 0.66 0.71 0.24 0.05 1107.32
@@
$\lambda_{1}\neq\lambda_{2}$
7pt $\mathbf{F}$ $\mathbf{F}$ 0 48.82 26.36 0.05 0.10 0.21 1.10 1.10 1144.53
7pt $\mathbf{F}$ $\mathbf{F}\lambda_{1}\lambda_{2}$ 0.9-0.9 23.66 11.66 0.55 0.61 0.67 0.26 0.06 1124.89
@@
Minimal Non-Minimal Sample AVG $(^{\circ})$ $\downarrow$ MED $(^{\circ})$ $\downarrow$ AUC@5 $\uparrow$ @10 @20 AVG $\epsilon(\lambda)$ $\downarrow$ MED $\epsilon(\lambda)$ $\downarrow$ Time (ms) $\downarrow$
$\lambda_{1}=\lambda_{2}$
7pt $\mathbf{F}$ 7pt $\mathbf{F}$ 0 13.11 5.72 0.21 0.38 0.55 0.86 0.85 190.04
@@
$\lambda_{1}\neq\lambda_{2}$
7pt $\mathbf{F}$ 7pt $\mathbf{F}$ 0 15.14 7.04 0.15 0.32 0.50 0.86 0.86 140.78
@@
Minimal Non-Minimal Sample AVG $(^{\circ})$ $\downarrow$ MED $(^{\circ})$ $\downarrow$ AUC@5 $\uparrow$ @10 @20 AVG $\epsilon(\lambda)$ $\downarrow$ MED $\epsilon(\lambda)$ $\downarrow$ Time (ms) $\downarrow$
$\lambda_{1}=\lambda_{2}$
7pt $\mathbf{F}$ 7pt $\mathbf{F}$ 0 30.22 22.80 0.05 0.11 0.23 0.85 0.84 170.37
@@
$\lambda_{1}\neq\lambda_{2}$
7pt $\mathbf{F}$ 7pt $\mathbf{F}$ 0 31.36 25.39 0.03 0.09 0.20 0.86 0.86 141.03
@@
Minimal Non-Minimal Sample AVG $(^{\circ})$ $\downarrow$ MED $(^{\circ})$ $\downarrow$ AUC@5 $\uparrow$ @10 @20 AVG $\epsilon(\lambda)$ $\downarrow$ MED $\epsilon(\lambda)$ $\downarrow$ Time (ms) $\downarrow$
$\lambda_{1}=\lambda_{2}$
7pt $\mathbf{F}$ 7pt $\mathbf{F}$ 0 30.74 22.21 0.06 0.15 0.26 0.87 0.87 180.49
@@
$\lambda_{1}\neq\lambda_{2}$
7pt $\mathbf{F}$ 7pt $\mathbf{F}$ 0 31.60 23.63 0.04 0.11 0.23 0.86 0.87 160.39
@@
Minimal Non-Minimal Sample AVG $(^{\circ})$ $\downarrow$ MED $(^{\circ})$ $\downarrow$ AUC@5 $\uparrow$ @10 @20 AVG $\epsilon(\lambda)$ $\downarrow$ MED $\epsilon(\lambda)$ $\downarrow$ Time (ms) $\downarrow$
$\lambda_{1}=\lambda_{2}$
7pt $\mathbf{F}$ 7pt $\mathbf{F}$ 0 17.24 6.45 0.21 0.36 0.51 0.86 0.85 160.49
@@
$\lambda_{1}\neq\lambda_{2}$
7pt $\mathbf{F}$ 7pt $\mathbf{F}$ 0 18.78 8.04 0.16 0.31 0.47 0.86 0.86 161.19
@@
Minimal Non-Minimal Sample AVG $(^{\circ})$ $\downarrow$ MED $(^{\circ})$ $\downarrow$ AUC@5 $\uparrow$ @10 @20 AVG $\epsilon(\lambda)$ $\downarrow$ MED $\epsilon(\lambda)$ $\downarrow$ Time (ms) $\downarrow$
$\lambda_{1}=\lambda_{2}$
7pt $\mathbf{F}$ 7pt $\mathbf{F}$ 0 29.09 21.46 0.05 0.12 0.24 0.86 0.84 150.03
@@
$\lambda_{1}\neq\lambda_{2}$
7pt $\mathbf{F}$ 7pt $\mathbf{F}$ 0 31.74 24.68 0.03 0.09 0.20 0.86 0.85 150.01
@@
Minimal Non-Minimal Sample AVG $(^{\circ})$ $\downarrow$ MED $(^{\circ})$ $\downarrow$ AUC@5 $\uparrow$ @10 @20 AVG $\epsilon(\lambda)$ $\downarrow$ MED $\epsilon(\lambda)$ $\downarrow$ Time (ms) $\downarrow$
$\lambda_{1}=\lambda_{2}$
7pt $\mathbf{F}$ 7pt $\mathbf{F}$ 0 36.74 31.81 0.05 0.10 0.19 0.86 0.85 140.01
@@
$\lambda_{1}\neq\lambda_{2}$
7pt $\mathbf{F}$ 7pt $\mathbf{F}$ 0 39.80 35.68 0.02 0.06 0.13 0.86 0.86 140.71
@@
Minimal Non-Minimal Sample AVG $(^{\circ})$ $\downarrow$ MED $(^{\circ})$ $\downarrow$ AUC@5 $\uparrow$ @10 @20 AVG $\epsilon(\lambda)$ $\downarrow$ MED $\epsilon(\lambda)$ $\downarrow$ Time (ms) $\downarrow$
$\lambda_{1}=\lambda_{2}$
7pt $\mathbf{F}$ 7pt $\mathbf{F}$ 0 25.16 16.80 0.08 0.17 0.31 0.88 0.88 1120.22
@@
$\lambda_{1}\neq\lambda_{2}$
7pt $\mathbf{F}$ 7pt $\mathbf{F}$ 0 32.47 24.35 0.04 0.10 0.21 0.86 0.84 1203.38
@@
Minimal Non-Minimal Sample AVG $(^{\circ})$ $\downarrow$ MED $(^{\circ})$ $\downarrow$ AUC@5 $\uparrow$ @10 @20 AVG $\epsilon(\lambda)$ $\downarrow$ MED $\epsilon(\lambda)$ $\downarrow$ Time (ms) $\downarrow$
$\lambda_{1}=\lambda_{2}$
7pt $\mathbf{F}$ 9pt $\mathbf{F}\lambda$ 1.5-1.5 15.85 2.43 0.38 0.52 0.64 0.13 0.08 1170.05
7pt $\mathbf{F}$ 12pt $\mathbf{F}\lambda_{1}\lambda_{2}$ 1.5-1.5 16.06 2.80 0.34 0.51 0.63 0.09 0.09 1180.12
@@
Minimal Non-Minimal Sample AVG $(^{\circ})$ $\downarrow$ MED $(^{\circ})$ $\downarrow$ AUC@5 $\uparrow$ @10 @20 AVG $\epsilon(\lambda)$ $\downarrow$ MED $\epsilon(\lambda)$ $\downarrow$ Time (ms) $\downarrow$
$\lambda_{1}=\lambda_{2}$
7pt $\mathbf{F}$ 9pt $\mathbf{F}\lambda$ 0.7-0.7 16.64 2.58 0.37 0.51 0.62 0.17 0.08 1205.99
7pt $\mathbf{F}$ 12pt $\mathbf{F}\lambda_{1}\lambda_{2}$ 0.7-0.7 15.98 3.11 0.32 0.49 0.62 0.10 0.09 1220.47